Which of the following assumptions implies that saving and investment are equal for a closed economy? a. Private saving is equal to zero. b. Public saving is equal to zero. c. The economy’s government is running neither a surplus nor a deficit. d. No assumption is necessary; saving and investment are equal for all closed economies.
Answer #1

Answer

The correct answer is (d) No assumption is necessary; saving and investment are equal for all closed economies.

At equilibrium Y = AE

When economy is closed AE = C + I + G

where C = consumption , I = investment and G = Government purchases

Hence As economy is in equilibrium implies Y = AE = C + I + G

=> I = Investment = Y - C - G

Total Saving(S) = Private saving + public saving

Private saving = Y - C - T where T = Taxes

Public Saving = T - G

Hence Total Saving(S) = Y - C - T + T - G = Y - C - G

Hence Total saving = Investment = Y - C - G in the economy.

This implies that no condition is required in order for saving and investment to be equal in a closed economy .

Hence, the correct answer is (d) No assumption is necessary; saving and investment are equal for all closed economies.